OpenStudy (anonymous):

COuld someone please help! im tired and i want to go to bed! ill give medals and whatever you need! A function is shown below: f(x) = x3 + 3x^2 - x - 3 Part A: What are the factors of f(x)? Show your work. Part B: What are the zeros of f(x)? Show your work. Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x).

OpenStudy (anonymous):

@dan815 @nincompoop please help im really sorry for bothering you but i need help

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@Australopithecus

OpenStudy (anonymous):

i am so sorry for tagging everyone but i really need help

OpenStudy (anonymous):

@paki

OpenStudy (dan815):

-1 is a root because if u plug in -1 it goes to 0

OpenStudy (anonymous):

so it could be a factor of f(x)

OpenStudy (anonymous):

i am so sorry if i come off as really stupid i am pretty much fried from all the work iv been doing

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

im so sorry for bothering you my brother just actually came home and helped me do it. i am really sorry for being a nuisance

OpenStudy (dan815):

okay soq

OpenStudy (dan815):

listen

OpenStudy (dan815):

a root means that there is a point when x= that root y=0

OpenStudy (dan815):

that means this polynomail must have a (x-r) (ax^2+bx+c) if r is the root term

OpenStudy (dan815):

now we say that -1 is a root so, (x+1) is a factor of x3 + 3x^2 - x - 3 to see what the (ax^2+bx+c) part what we do is x3 + 3x^2 - x - 3 ----------------- = use syntetic division and see what this is (x+1) this must be the ax^2+bx+c part except you would have solved for a b and c part Then you can see if you can continue to factor this (ax^2+bx+c) and make (x-r) (ax^2+bx+c) = (x-r1) (x-r2)(x-r3) into this completely product ant root form, r1,r2,r3 are the roots in this case r1=-1 (x-(-1))(x-r2)(x-r3) = (x+1)(x-r2)(x-r3) you still have to find r2 and r3 not that isnt a square or cube on r2 and r3 that is an index.